OpenStudy (anonymous):

how to get cos(-3pie/4)

OpenStudy (luigi0210):

Uhm, you get it by going 135 degrees clockwise on the unit circle?

OpenStudy (anonymous):

is there another way?

OpenStudy (anonymous):

 \[\cos -3\pi/4 = -\sqrt{2}/2\]